INAF queue observing, 2018 June 25-26

Observer: A. Rossi
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn (on call)
Instrument: LUCI

Summary

We gave 2 hrs at the beginning of the night to LBTO to solve AGW problems: loosing on large offsets, some software problem At the sunset the sky is clear
instruments: MODS2 red grating is not working well
Night duration (12\xB0 el twilight) 3:35- 11:10 UT = 7 h 35m

Important : guiding with LUCI1 needs to abort the pause for collimation when the TiO says collimation is ok.
  • This happens for every active preset, i.e. beginning of a script. You do this on telsvc gui

Observing time: h m (%)

Weather time loss: h m (%)

technical time loss * *(9%) but if we count for 50% LUCI1 images useless, it is another 10%

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
STD BRI2202 LUCI1+2 ima,H2,Ks     per Moscadelli
Moscadelli G01268 LUCI1+2 ima,H2,Ks     we toke more images, because 50% of LUCI1 and 10% on LUCI2 look not well collimated, show elongated or double stars, check but should be completed
Caccianiga GB6J001307 +205335 MODS1 red grating 2400/3600 mods1r.20180626.0005-0008 completed PI confirmed
STD Feige110 MODS1 red grating   mods1r.20180626.0014-0016 per Caccianiga
CALIBRATIONS MODS1 red grating Caccianiga
CALIBRATIONS MODS1 dualgrating imaging flats, maybe useful
CALIBRATIONS Dome Flats LUCI1+2 H2, Ks Per Moscadelli

Details

3:00-5:00 UT LBTO work on AGW3

5:00 UT Configure LUCI1+2 We start with Moscadelli program

5:10 UT Slew to standard for Msocadelli. FS140 is too close to the Moon (6 deg, saturates). Use FS149. TO does collimation.

5:17 UT send FS149_H2_bino DO NOT USE THIS STANDARD BUT THEONE AT THE END!
  • seeing 0.7" on guider
  • even without tracking looks not always fine, especially in LUCI1. Shows elongations. enough for calibration.
  • luci1.0014-18 , luci2.0013-17
  • 5:28 UT send FS149_K_bino
  • same as above
  • but I plan to take another one at end of the night.
  • luci1.0019-23 , luci2.0018-22

5:32 UT slew to G01268 Moscadelli program
  • first H2 filter
  • luci1.0024-28 , luci2.0023-27
  • seeing 0.7" on guider
  • on luci 1 stars look elongated, a bit on luci2 too. STart a script with half NDIT, double offsets.
  • luci1.0029 , luci2.0028 aborted wrong preset
  • luci1.0030-42 , luci2.0029-43
  • seeing 0.8" on guider
  • luci2 ok, luci1 still shows nord-sud elongation, or double stars sometimes
  • seeing 0.6" on guider
  • last 2 luci1 images much better luci1.0040-42
  • 7:01 lost preset * 5 min loss*
  • sent again from where we stopped
  • luci1.0043-48 , luci2.0044-47
  • luci1.0046 and so on is better, maybe useful? check with PI
  • 7:21 lost preset * 20 min loss*
  • 7:40 send again from where we stop
  • luci1.0049-54 , luci2.0048-53
  • LUCI1 ok from luci1.0050
  • seeing 1.0" on guider at 7:46

* Ks filter
  • 8:01
  • luci1.0055-64 , luci2.0054-63 luci1.55 triangle stars, after 56 are ok, 63 64 are bad
  • LUCI2 59-63 ecc are bad
  • send again need 4 offsets more with good images
  • luci1.0065-71 , luci2.0064-70 nothing good on boths sides!
  • send again but with fainter guiding star
  • picked diffenrent guiding stars due to crowded field, luci1 may be few arcsec offset
  • luci1.0072-81 , luci2.0071-80
  • seeing 1.1" on guider at 8:30
  • luci2 is ok now, maybe fainter guider helped, maybe reset collimaiton , who knows
  • just luci2 74 and 75 are bad, luci 1 always elongated maybe ok
  • 8:35 * 5 min loss* tried to reset optics, then continue from where we stop
  • seeing 1.2" on guider at 8:40
  • now they look fine.finished with Ks filter

8:48 send H2 filter again
  • need other 10 H2 filter at least to get good enough exposures (48 offsets summing both sides to get 2 hrs recquested, with this NDIT halved )
  • luci1.0082-88 , luci2.0080-87
  • seeing 0.9" on guider at 8:53
  • we are very low, seeing jumps between 1.1 and1.4 after image luci1.85 and luci2.84
  • got enough good images, in total we should have 2 hours on H2filter

9:15 UT slew to standard BRI2202
  • pointing problem, restart GCS on luci1 5 min loss
  • first H2 filter
  • luci1.0089-93 , luci2.0088-92
  • 9:30 Ks filter
  • luci1.0094- , luci2.0093-

9:36 switch to MODS

9:56 UT slew to Caccianiga GB6J001307 +205335
  • seeing 1.1" on guider
  • acq mods1r.20180626.0002-0004
  • 10:08 spec mods1r.20180626.0005-0008
  • seeing 0.9" on guider, clear continuum in the spectra
  • run again script at 10:46, took 1 exposure, we move to std
  • seeing 0.8" on guider at 10:50

10:59 Slew to STD Feige110
  • acq mods1r.20180626.0009-0013
  • 11:13 spec mods1r.20180626.0014-0016

MODS1 CALIBRATIONS for Caccianiga
  • ==mods1r.20180626.0017-0022==slitflats 1.2 red grating
  • mods1r.20180626.0023-0027 grpixflats red grating
  • mods1r.20180626.0029-0031 slitflats 5 red grating
  • mods1r.20180626.0032-0046 imflats dual grating, you never know

*LUCI1+2 CALIBRATIONS
  • luci1.0112-122 , luci2.0111-121 H2 flats, first 5 are darks to discharge
  • luci1.0123-131 , luci2.0122-130 Ks flats , first 5 are darks to discharge

-- %USERSIG{AndreaRossi - 2018-06-26}%

Comments

 
Topic revision: r5 - 30 Jun 2018, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback